### v2

# Antisymmetric SU(2) adjoint generator as the universal origin of geometric phases in quantum two-level systems and classical polarization

Abstract
We establish that the antisymmetric part of the adjoint SU(2) generator serves as the universal algebraic origin of geometric phases across classical polarization optics and quantum two-level systems. For classical ideal retarders, we demonstrate that the antisymmetric block of the Mueller matrix exclusively determines the geometric phase, encoding the angular-velocity pseudovector governing Stokes-vector evolution on the Poincaré sphere. Remarkably, the symmetric component is geometrically neutral. Quantum mechanically, the identical antisymmetric generator emerges in the adjoint SU(2) action on the Bloch sphere and fully governs the geometric phase of pure qubit states, independent of adiabaticity or cyclicity. This unified framework provides direct operational access to geometric phases via measured Mueller matrices or quantum process tomography, enabling precise control and diagnosis of geometric effects in both classical and quantum domains.
I Introduction
Geometric phases appear in a wide variety of wave phenomena, from classical polarization optics to quantum two-level dynamics (see, e.g., Ref. [1] for a recent review). In classical optics, Pancharatnam’s seminal work on interference between different polarization states introduced a geometric connection on the space of polarization states and led to what is now known as the Pancharatnam–Berry (PB) phase [2, 6]. In quantum mechanics, Berry showed that a quantum system undergoing adiabatic and cyclic evolution acquires a geometric phase that depends solely on the path traced in parameter space [3]. For two-level systems, these classical and quantum viewpoints share the same SU(2) and Bloch/Poincaré-sphere structure. This concept was subsequently generalized to nonadiabatic and noncyclic evolutions by Aharonov and Anandan [4] and by Samuel and Bhandari [5]. For polarized light, a comprehensive review of geometric and topological phase phenomena is given in Ref. [6]. Further generalizations to mixed quantum states in an interferometric setting were developed by Sjöqvist et al. [7].
Although these developments are often presented in different formalisms (Jones and Mueller calculus in classical polarization, and Hilbert-space or SU(2)/Bloch-sphere representations in quantum mechanics) they share the same geometric origin: a curve on a space of states that can be identified with the two-sphere $S^{2}$. In both cases, pure states correspond to points on the Poincaré or Bloch sphere, and physically relevant evolutions are represented by rotations in $\mathrm{SO}(3)$ induced by the adjoint action of $\mathrm{SU}(2)$.
Here we make this correspondence explicit at the algebraic level and reveal a fundamental structure: the antisymmetric part of the adjoint $\mathrm{SU}(2)$ generator provides a universal kernel for geometric phases in both classical polarization optics and quantum two-level systems. This perspective not only unifies these seemingly disparate phenomena but also provides direct operational access to geometric phases from experimental measurements. On the classical side, we consider arbitrary ideal retarders described by Mueller matrices and show that their geometric action is completely encoded in the antisymmetric $3\times 3$ block acting on the Stokes vector. This antisymmetric block encodes the angular-velocity pseudovector that governs the tangential component of the Stokes-vector motion on the Poincaré sphere and therefore fully determines the associated geometric phase. The complementary symmetric block may reshape the mapping between Stokes vectors, but it is geometrically neutral in the sense that it does not contribute to the geometric phase.
On the quantum side, we show that the same antisymmetric structure appears in the adjoint action of any $\mathrm{SU}(2)$ unitary operator on a qubit. The Bloch vector of a pure qubit state evolves as $\mathbf{r}(t)=R(t)\,\mathbf{r}(0)$, where $R(t)\in\mathrm{SO}(3)$ is a continuous one-parameter family of rotations and its instantaneous generator $A(t)$, defined by $\dot{R}(t)=A(t)R(t)$, is an antisymmetric matrix. We demonstrate that the geometric phase of a pure qubit state is fully determined by this antisymmetric generator, independently of whether the evolution is adiabatic or nonadiabatic, cyclic or noncyclic. The role of the antisymmetric generator is again to encode the instantaneous angular velocity of the Bloch vector and, more specifically, the tangential component of this velocity on the Bloch sphere, which is the only part that contributes to the geometric phase.
These results identify a single algebraic structure underlying PB phases in classical polarization and geometric phases in quantum two-level systems: the antisymmetric part of the adjoint $\mathrm{SU}(2)$ generator. This viewpoint is also operational. In classical experiments, the relevant antisymmetric block can be extracted directly from a measured Mueller matrix of an ideal retarder. In quantum experiments, the same antisymmetric generator can be reconstructed from process tomography through the adjoint map acting on the Bloch vector. In both cases this yields a simple criterion for isolating geometric-phase contributions and for diagnosing transformations where geometric phases are controlled or suppressed.
The remainder of the paper is organized as follows. In Sec. II we analyze the Mueller matrix of an arbitrary ideal retarder and show that its antisymmetric block is the unique generator of geometric phase on the Poincaré sphere. In Sec. III we connect this antisymmetric generator with the tangential angular velocity of the Stokes vector and with the solid angle enclosed by its trajectory. Section IV extends the formalism to quantum two-level systems by studying the adjoint action of $\mathrm{SU}(2)$ on the Bloch sphere. In Sec. V we present illustrative classical and quantum examples that highlight the operational content of the theory. Section VI discusses possible extensions to more general quantum channels and higher-dimensional systems, and Sec. VII summarizes our conclusions.
II Classical polarization: Mueller retarders and antisymmetric generator
We first consider the classical case of an ideal retarder acting on a fully polarized beam. The retarder is described in the Jones formalism by a unitary $2\times 2$ operator of the form [8]
$U=\exp\!\left[-\frac{i}{2}\,\delta\,\mathbf{n}\cdot\bm{\sigma}\right],$ (1)
where $\delta$ is the retardance, $\mathbf{n}$ is a real unit vector that specifies the eigenpolarizations on the Poincaré sphere, and $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the Pauli matrices. The pair $(\delta,\mathbf{n})$ provides a complete parametrization of pure retarders: $\delta$ determines the rotation angle on the Poincaré sphere and $\mathbf{n}$ the rotation axis.
The Pauli-algebraic approach to polarization optics was developed by Whitney in her early work on Pauli-algebraic operators in polarization optics [8]. A systematic development of a vectorial, pure operatorial Pauli-algebraic formalism for device and state operators has been given by Tudor [9, 10], providing an alternative but closely related framework to the standard Jones and Mueller calculi for deterministic (also called nondepolarizing) systems.
The corresponding action on the Stokes vector is given by a real $3\times 3$ rotation matrix $R\in\mathrm{SO}(3)$ acting as
$\mathbf{s}_{\mathrm{out}}=R\,\mathbf{s}_{\mathrm{in}},$ (2)
where $\mathbf{s}=(S_{1},S_{2},S_{3})^{\mathsf{T}}$ denotes the normalized Stokes vector of a fully polarized state. The matrix $R$ is the adjoint $\mathrm{SO}(3)$ representation of the $\mathrm{SU}(2)$ operator $U$ and takes the standard Rodrigues form (see, e.g., Ref. [11, 12])
$R=\cos\delta\,\mathbb{I}_{3}+(1-\cos\delta)\,\mathbf{n}\mathbf{n}^{\mathsf{T}}+\sin\delta\,[\mathbf{n}]_{\times},$ (3)
where $\mathbb{I}_{3}$ is the $3\times 3$ identity matrix, $\mathbf{n}\mathbf{n}^{\mathsf{T}}$ is the rank-one projector onto $\mathbf{n}$, and $\mathbf{n}]_{\times}$ is the cross-product matrix
$\mathbf{n}]_{\times}=\begin{pmatrix}0&-n_{3}&n_{2}\\ n_{3}&0&-n_{1}\\ -n_{2}&n_{1}&0\end{pmatrix}$, (4)
which satisfies $\mathbf{n}]_{\times}\,\mathbf{v}=\mathbf{n}\times\mathbf{v}$ for any vector $\mathbf{v}$.
In the Mueller-matrix formalism, a pure (or ideal) retarder is represented by a $4\times 4$ Mueller matrix [13]:
$M=\begin{pmatrix}1&\mathbf{0}^{\mathsf{T}}\\ \mathbf{0}&R\end{pmatrix}$, (5)
where the upper-left unit element $m_{00}=1$ reflects no attenuation and the $3\times 3$ submatrix $R$ acts on the reduced Stokes vector. Thus, for ideal retarders, all polarimetric structure relevant to the geometric phase is contained in $R$.
To identify the geometric generator, we decompose $R$ into its symmetric and antisymmetric parts,
$R=S+A,\qquad S=\frac{1}{2}\left(R+R^{\mathsf{T}}\right),\quad A=\frac{1}{2}\left(R-R^{\mathsf{T}}\right),$ (6)
with $S^{\mathsf{T}}=S$ and $A^{\mathsf{T}}=-A$. Using Eq. (3) and the antisymmetry of $\mathbf{n}]_{\times}$, a straightforward calculation gives
$S=\cos\delta\,\mathbb{I}_{3}+(1-\cos\delta)\,\mathbf{n}\mathbf{n}^{\mathsf{T}},$ (7)
$A=\sin\delta\,[\mathbf{n}]_{\times}.$ (8)
The symmetric part $S$ thus depends only on $\delta$ and on the projector along $\mathbf{n}$, whereas the antisymmetric part $A$ is proportional to the cross-product matrix $\mathbf{n}]_{\times}$ with coefficient $\sin\delta$.
From a geometric viewpoint, the antisymmetric matrix $A=\sin\delta\,[\mathbf{n}]_{\times}$ can be identified with an axial vector $\bm{\Omega}=\sin\delta\,\mathbf{n}$ whose direction and sign fix the sense of rotation of Stokes vectors on the Poincaré sphere. Reversing $\bm{\Omega}$ reverses the circulation of any trajectory and flips the sign of the corresponding geometric phase. In this sense, the antisymmetric adjoint generator carries the chirality of the evolution on the sphere, whereas the symmetric part $S$ is achiral and geometrically neutral.
Equations (7) and (8) show that the action of a pure retarder on the Poincaré sphere can be decomposed into two algebraically distinct components: a symmetric mapping that deforms the distribution of Stokes vectors along and orthogonal to the eigenaxis $\mathbf{n}$, and an antisymmetric generator that produces a rotation around $\mathbf{n}$. As we will show in Sec. III, only the antisymmetric generator $A$ contributes to the geometric phase: it encodes the instantaneous angular velocity of the Stokes vector and, in particular, its tangential component on the sphere. The symmetric part $S$ may modify the detailed mapping of Stokes vectors, but it does not contribute to the accumulated geometric phase.
For any ideal retarder the antisymmetric $3\times 3$ block of the Mueller matrix is therefore $A=\sin\delta\,[\mathbf{n}]_{\times},$ and this matrix will be identified in the following sections as the universal classical kernel of geometric phase.
Figure 1: Representative closed trajectory on the Poincaré/Bloch sphere in $(S_{1},S_{2},S_{3})$ coordinates. The closed curve and its shaded interior represent the solid angle enclosed by the Stokes/Bloch vector. The instantaneous tangential component of the angular velocity, $\bm{\Omega}_{\perp}$, lies in the tangent plane to the sphere and is generated exclusively by the antisymmetric adjoint operator $A=\sin\delta\,[\mathbf{n}]_{\times}$ for classical pure retarders and by $A_{U}=\sin\delta\,[\mathbf{n}]_{\times}$ for quantum two-level unitaries. This common antisymmetric generator fully determines the geometric phase. The actual velocity of the Stokes/Bloch vector along the trajectory is $\dot{\mathbf{s}}=\bm{\Omega}_{\perp}\times\mathbf{s}$.
III Geometric phase and tangential angular velocity on the Poincaré sphere
The decomposition (6)–(8) shows that the polarimetric action of an ideal retarder on the Poincaré sphere is governed by an antisymmetric matrix $A$ of the form
$A=[\bm{\Omega}]_{\times},\qquad\bm{\Omega}=\sin\delta\,\mathbf{n},$ (9)
where we have introduced the angular-velocity vector $\bm{\Omega}$ associated with the rotation around $\mathbf{n}$ in the usual sense of rigid-body kinematics on the sphere. Indeed, for any Stokes vector $\mathbf{s}$ we have
$A\,\mathbf{s}=[\bm{\Omega}]_{\times}\,\mathbf{s}=\bm{\Omega}\times\mathbf{s},$ (10)
so that the infinitesimal action of the retarder on $\mathbf{s}$ can be written in the kinematic form
$\frac{d\mathbf{s}}{dt}=\bm{\Omega}(t)\times\mathbf{s}(t),$ (11)
where $t$ denotes any convenient parameter along the evolution (for instance, the physical time, the propagation distance inside the medium, or a control parameter such as the rotation angle of a wave plate). In this sense $\bm{\Omega}(t)$ represents the instantaneous angular velocity of $\mathbf{s}(t)$ on the Poincaré sphere, and Eq. (11) is the standard equation of motion for a unit vector driven by such an angular velocity.
More generally, an arbitrary retarder can be viewed as a concatenation of elementary retarders with piecewise constant parameters $(\delta,\mathbf{n})$ or, in the continuum limit, as generated by a smooth angular-velocity field $\bm{\Omega}(t)$ along the evolution parameter. The net action is always equivalent to a single rotation on the sphere with some effective generator $\bm{\Omega}_{\mathrm{eff}}]_{\times}$, although $\bm{\Omega}_{\mathrm{eff}}$ is in general a nontrivial function of the full history of $\bm{\Omega}(t)$ due to the noncommutativity of rotations. A physically relevant example of such a continuum limit is provided by an optical fiber whose local birefringence and eigenpolarizations vary smoothly along the propagation direction, so that the Stokes vector experiences a continuously varying angular velocity on the Poincaré sphere.
The geometric phase associated with the evolution of a fully polarized state can be understood as a property of the trajectory $\mathbf{s}(t)$ on the Poincaré sphere, rather than as a property of the underlying Jones vector itself. For a closed loop $C$ traced by $\mathbf{s}(t)$, the Pancharatnam–Berry (PB) phase is [2, 3, 5, 6]
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\,\Omega_{\mathrm{solid}}(C),$ (12)
where $\Omega_{\mathrm{solid}}(C)$ is the oriented solid angle subtended by $C$ at the center of the sphere (see Fig. 1). The expression (12) can be derived from the Pancharatnam connection $\gamma=\arg\langle\psi(0)|\psi(T)\rangle-\text{(dynamical phase)}$ by recognizing that the overlap between initial and final states on the sphere is governed by the solid angle enclosed by the trajectory. For a comprehensive geometric treatment see [6, 5].
This expression can be written in differential form as a line integral over the curve or, equivalently, as a surface integral over any surface $\Sigma$ bounded by $C$.
Figure 2: Local geometry of the Stokes/Bloch vector and the antisymmetric adjoint generator at a point $\mathbf{s}$ on the Poincaré/Bloch sphere. For visual clarity the point is shown at the north pole of the sphere, but the construction is local and applies to any point. The unit vector $\mathbf{s}$ is normal to the sphere and to the tangent plane $T_{\mathbf{s}}$ at that point. The instantaneous angular velocity $\bm{\Omega}$ is decomposed into a component $\bm{\Omega}_{\parallel}$ parallel to $\mathbf{s}$ and a component $\bm{\Omega}_{\perp}$ that lies in the tangent plane. The velocity of the Stokes/Bloch vector is $\dot{\mathbf{s}}=\bm{\Omega}_{\perp}\times\mathbf{s}$; it is therefore tangent to the sphere and orthogonal to $\bm{\Omega}_{\perp}$ within the tangent plane. Reversing $\bm{\Omega}_{\perp}$ reverses the circulation of the trajectory and the sign of the geometric phase.
To make the connection with Eq. (11), we recall that the solid angle can be expressed as a surface integral of the Berry curvature on the sphere. Let $\mathbf{s}(\theta,\phi)$ denote the unit vector corresponding to spherical angles $(\theta,\phi)$, with $0\leq\theta\leq\pi$ and $0\leq\phi<2\pi$, and let $\Sigma$ be a surface on $S^{2}$ with boundary $C$. Then
$\Omega_{\mathrm{solid}}(C)=\iint_{\Sigma}\mathcal{F}(\theta,\phi)\,d\theta\,d\phi,$ (13)
where the Berry curvature two-form on the sphere is (see, e.g., Refs. [3, 14])
$\mathcal{F}(\theta,\phi)=\mathbf{s}\cdot\left(\frac{\partial\mathbf{s}}{\partial\theta}\times\frac{\partial\mathbf{s}}{\partial\phi}\right)$ (14)
and reduces to $\mathcal{F}=\sin\theta$ in the usual parametrization. Equation (12) then becomes
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\iint_{\Sigma}\mathbf{s}\cdot\left(\frac{\partial\mathbf{s}}{\partial\theta}\times\frac{\partial\mathbf{s}}{\partial\phi}\right)d\theta\,d\phi.$ (15)
On the other hand, the kinematic equation (11) implies that the instantaneous motion of $\mathbf{s}(t)$ on the sphere is characterized by the angular velocity $\bm{\Omega}(t)$. We can decompose $\bm{\Omega}(t)$ into components parallel and orthogonal to $\mathbf{s}(t)$,
$\bm{\Omega}(t)=\bm{\Omega}_{\parallel}(t)+\bm{\Omega}_{\perp}(t),\qquad\bm{\Omega}_{\parallel}=(\bm{\Omega}\cdot\mathbf{s})\,\mathbf{s},\quad\bm{\Omega}_{\perp}=\bm{\Omega}-\bm{\Omega}_{\parallel},$ (16)
where $\bm{\Omega}_{\parallel}$ is the component parallel to $\mathbf{s}$ and $\bm{\Omega}_{\perp}$ lies in the tangent plane of the sphere at $\mathbf{s}$.
The local geometry of these quantities is illustrated in Fig. 2. At a given time $t$, the unit Stokes/Bloch vector $\mathbf{s}(t)$ is normal to the sphere and therefore normal to the tangent plane $T_{\mathbf{s}}$ at that point. The instantaneous angular velocity vector $\bm{\Omega}(t)$ can be decomposed into a component parallel to $\mathbf{s}(t)$ and a component tangent to the sphere, $\bm{\Omega}(t)=\bm{\Omega}_{\parallel}(t)+\bm{\Omega}_{\perp}(t)$, with $\bm{\Omega}_{\parallel}(t)=(\bm{\Omega}\cdot\mathbf{s})\,\mathbf{s}$ and $\bm{\Omega}_{\perp}(t)=\bm{\Omega}(t)-\bm{\Omega}_{\parallel}(t)$. Here $\bm{\Omega}_{\parallel}$ is collinear with $\mathbf{s}$ and does not affect the motion on the sphere, whereas $\bm{\Omega}_{\perp}$ lies entirely in the tangent plane $T_{\mathbf{s}}$.
The kinematic equation $\dot{\mathbf{s}}=\bm{\Omega}\times\mathbf{s}$ shows that the velocity $\dot{\mathbf{s}}(t)$ of the Stokes/Bloch vector is also tangent to the sphere and is in fact determined solely by the tangential component of the angular velocity, $\dot{\mathbf{s}}(t)=\bm{\Omega}_{\perp}(t)\times\mathbf{s}(t)$, since $\bm{\Omega}_{\parallel}(t)\times\mathbf{s}(t)=\mathbf{0}$. As a consequence, both $\bm{\Omega}_{\perp}(t)$ and $\dot{\mathbf{s}}(t)$ lie in the tangent plane and are mutually orthogonal there. Reversing $\bm{\Omega}_{\perp}$ reverses the circulation of the trajectory on the sphere and flips the sign of the associated geometric phase. This makes explicit that the tangential component of the antisymmetric adjoint generator encodes not only the magnitude of the instantaneous angular velocity, but also the handedness of the evolution on the Poincaré/Bloch sphere.
Equation (11) shows that only $\bm{\Omega}_{\perp}$ contributes to the motion of $\mathbf{s}$, since
$\bm{\Omega}_{\parallel}(t)\times\mathbf{s}(t)=\mathbf{0},\qquad\bm{\Omega}_{\perp}(t)\times\mathbf{s}(t)=\bm{\Omega}(t)\times\mathbf{s}(t).$ (17)
This decomposition reveals why only the antisymmetric generator contributes to geometric phases: $\bm{\Omega}_{\parallel}$ is collinear with $\mathbf{s}$ and does not induce any motion on the sphere, whereas $\bm{\Omega}_{\perp}$ fully drives the evolution of $\mathbf{s}(t)$. In particular, the component of $\bm{\Omega}_{\perp}$ along the local normal to the osculating plane determines the geodesic curvature and thus the sign and magnitude of the geometric phase.
The PB phase can be expressed in terms of $\bm{\Omega}_{\perp}(t)$ by noting that the solid angle enclosed by the trajectory can be written as an integral over time of the oriented area swept by $\mathbf{s}(t)$ on the sphere. A convenient way to express this is to parametrize the curve by $t$ and write
$\Omega_{\mathrm{solid}}(C)=\int_{0}^{T}dt\,\kappa_{\mathrm{g}}(t)\,v(t),$ (18)
where $v(t)=|\dot{\mathbf{s}}(t)|$ is the speed of $\mathbf{s}(t)$ along the curve (with respect to the standard metric on $S^{2}$), and $\kappa_{\mathrm{g}}(t)$ is the geodesic curvature of the trajectory. For unit-speed motion one has the classical relation $\kappa_{\mathrm{g}}=\bm{\Omega}_{\perp}\cdot\mathbf{u}_{\mathrm{n}}$, where $\mathbf{u}_{\mathrm{n}}$ is the normal to the osculating plane, and the sign of $\kappa_{\mathrm{g}}$ captures the orientation of the curve. Thus the geometric phase can be written in the kinematic form
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\int_{0}^{T}dt\,\kappa_{\mathrm{g}}(t)\,v(t),$ (19)
which shows explicitly that $\gamma_{\mathrm{geom}}$ depends only on the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity.
Equations (9)–(19) establish that, for ideal retarders, the geometric phase is fully determined by the antisymmetric generator $A$. The symmetric part $S$ of the Mueller matrix does not enter in Eq. (11), and therefore it does not contribute to the solid angle or to $\gamma_{\mathrm{geom}}$. In other words, the antisymmetric $3\times 3$ block $A$ of the Mueller matrix contains the complete geometric content of the retarder: it specifies the instantaneous angular velocity $\bm{\Omega}$ and, in particular, its tangential component $\bm{\Omega}_{\perp}$ that controls the PB phase through Eq. (19).
Although the discussion above is phrased in terms of classical polarization, the kinematic equation (11) and the decomposition (16) are purely geometric statements about curves on the two-sphere. As such, they apply equally well to the evolution of Bloch vectors representing pure states of a quantum two-level system. In the next section we show that the adjoint action of any $\mathrm{SU}(2)$ unitary operator on the Bloch sphere is generated by an antisymmetric matrix of the same form (8), and that the corresponding geometric phase is again fully determined by this antisymmetric generator.
IV Quantum two-level systems: adjoint SU(2) action and Bloch-sphere geometric phase
We now turn to quantum two-level systems and show that the same antisymmetric generator that appears in the Mueller description of pure retarders also governs geometric phases on the Bloch sphere. Any qubit state can be represented by a Bloch vector $\mathbf{r}$, with pure states lying on the unit sphere and mixed states in its interior, and their dynamics is generated by $\mathrm{SU}(2)$ unitaries in the adjoint representation [12, 15]. For cyclic evolutions of pure states, the Aharonov–Anandan phase [4] can be written as
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\,\Omega_{\mathrm{solid}}(C),$ (20)
in direct analogy with the classical PB phase.
Consider a qubit with density matrix
$\rho=\frac{1}{2}\left(\mathbb{I}_{2}+\mathbf{r}\cdot\bm{\sigma}\right),$ (21)
where $\mathbf{r}\in\mathbb{R}^{3}$ is the Bloch vector, $|\mathbf{r}|\leq 1$, and $\bm{\sigma}$ are the Pauli matrices. Pure states correspond to $|\mathbf{r}|=1$, so that $\mathbf{r}$ lies on the Bloch sphere.
Let $U(t)\in\mathrm{SU}(2)$ be a unitary evolution operator acting on the qubit, with $U(0)=\mathbb{I}_{2}$. The density matrix evolves as
$\rho(t)=U(t)\,\rho(0)\,U^{\dagger}(t),$ (22)
which induces an adjoint action on the Bloch vector,
$\mathbf{r}(t)=R_{U}(t)\,\mathbf{r}(0),$ (23)
where $R_{U}(t)\in\mathrm{SO}(3)$ is a real rotation matrix. The mapping $U\mapsto R_{U}$ is the adjoint representation of $\mathrm{SU}(2)$ on $\mathbb{R}^{3}$ and is formally identical to the mapping that relates the Jones operator (1) to the Poincaré-sphere rotation (3).
Any $U(t)\in\mathrm{SU}(2)$ can be written as
$U(t)=\exp\!\left[-\frac{i}{2}\,\delta(t)\,\mathbf{n}(t)\cdot\bm{\sigma}\right],$ (24)
where $\delta(t)$ is a time-dependent rotation angle and $\mathbf{n}(t)$ is a unit vector specifying the instantaneous rotation axis in Bloch space. Inserting Eq. (24) into the adjoint map and using standard $\mathrm{SU}(2)\to\mathrm{SO}(3)$ identities [12], one finds that $R_{U}(t)$ has exactly the same structure as Eq. (3),
$R_{U}(t)=\cos\delta(t)\,\mathbb{I}_{3}+\bigl[1-\cos\delta(t)\bigr]\,\mathbf{n}(t)\mathbf{n}^{\mathsf{T}}(t)+\sin\delta(t)\,[\mathbf{n}(t)]_{\times}.$ (25)
The corresponding antisymmetric and symmetric parts are
$S_{U}(t)=\cos\delta(t)\,\mathbb{I}_{3}+\bigl[1-\cos\delta(t)\bigr]\,\mathbf{n}(t)\mathbf{n}^{\mathsf{T}}(t),$ (26)
$A_{U}(t)=\sin\delta(t)\,[\mathbf{n}(t)]_{\times}.$ (27)
Thus, the adjoint action of $U(t)$ on the Bloch vector is generated by an antisymmetric matrix of the form
$A_{U}(t)=[\bm{\Omega}(t)]_{\times},\qquad\bm{\Omega}(t)=\sin\delta(t)\,\mathbf{n}(t),$ (28)
which leads to the kinematic equation
$\frac{d\mathbf{r}(t)}{dt}=\bm{\Omega}(t)\times\mathbf{r}(t),$ (29)
identical in form to Eq. (11). In this picture $\mathbf{r}(t)$ is a unit vector on the Bloch sphere and $\bm{\Omega}(t)$ plays the role of its instantaneous angular velocity on the sphere, now determined by the underlying $\mathrm{SU}(2)$ Hamiltonian, in direct analogy with the Stokes-vector evolution discussed in Sec. III.
As in the classical retarder case, the overall unitary evolution over a finite interval can always be represented by a single effective rotation on the Bloch sphere with generator $\bm{\Omega}_{\mathrm{eff}}]_{\times}$, although $\bm{\Omega}_{\mathrm{eff}}$ is in general a nontrivial functional of the full history of $\bm{\Omega}(t)$ because of the noncommutativity of $\mathrm{SU}(2)$ rotations.
As in the classical case, we can decompose $\bm{\Omega}(t)$ into components parallel and orthogonal to $\mathbf{r}(t)$,
$\bm{\Omega}(t)=\bm{\Omega}_{\parallel}(t)+\bm{\Omega}_{\perp}(t),$ (30)
with $\bm{\Omega}_{\parallel}=(\bm{\Omega}\cdot\mathbf{r})\,\mathbf{r}$ and $\bm{\Omega}_{\perp}=\bm{\Omega}-\bm{\Omega}_{\parallel}$.
The geometric phase associated with the evolution of a pure qubit state can again be written as a property of the trajectory $\mathbf{r}(t)$ on the Bloch sphere. For a cyclic evolution where $\mathbf{r}(T)=\mathbf{r}(0)$, the Aharonov–Anandan geometric phase [4] is given by
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\,\Omega_{\mathrm{solid}}(C),$ (31)
where $C$ is the closed Bloch trajectory and $\Omega_{\mathrm{solid}}(C)$ is the corresponding solid angle. This expression holds independently of whether the evolution is adiabatic or nonadiabatic and reduces to Berry’s original formula in the adiabatic case.
By the same geometric arguments as in Sec. III, the solid angle can be expressed in terms of the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity of $\mathbf{r}(t)$ on the sphere. In particular, one can write
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\int_{0}^{T}dt\,\kappa_{\mathrm{g}}(t)\,v(t),$ (32)
where $v(t)=|\dot{\mathbf{r}}(t)|$ is the speed of the Bloch vector along $C$ and $\kappa_{\mathrm{g}}(t)$ is the geodesic curvature of the trajectory. The geodesic curvature depends only on the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity and vanishes for geodesic motion on the sphere. Equation (32) therefore shows that the geometric phase of a pure qubit state is entirely determined by the antisymmetric generator $A_{U}(t)$.
From the operational point of view, this means that the geometric content of a unitary qubit evolution can be extracted solely from the antisymmetric part of the adjoint map $R_{U}(t)$. In experiments where the quantum process is characterized via qubit process tomography, one reconstructs the completely positive trace-preserving map $\mathcal{E}$ on qubit density matrices $\rho$, i.e., the map $\rho\mapsto\mathcal{E}(\rho)$, and the induced action on the Bloch vector,
$\mathbf{r}_{\mathrm{out}}=R_{\mathcal{E}}\,\mathbf{r}_{\mathrm{in}}+\mathbf{t},$ (33)
where $R_{\mathcal{E}}$ is a real $3\times 3$ matrix and $\mathbf{t}$ is a translation vector.
V Illustrative examples
In this section we present two simple examples that illustrate the role of the antisymmetric adjoint generator as the unique source of geometric phase. The first is a classical polarization experiment based on a rotating wave plate, and the second is a quantum spin-$1/2$ system in a cyclically varying magnetic field.
V.1 Classical example: rotating wave plate on the Poincaré sphere
Consider a lossless ideal wave plate with retardance $\delta$ and fast axis defined by a unit vector $\mathbf{n}(\alpha)$ on the Poincaré sphere, where $\alpha$ denotes a rotation angle of the plate around the propagation direction. For definiteness, we take the fast axis to lie initially in the equatorial plane and to rotate around the $S_{3}$ axis, so that
$\mathbf{n}(\alpha)=\bigl(\cos 2\alpha,\;\sin 2\alpha,\;0\bigr)^{\mathsf{T}}.$ (34)
The Jones operator of the plate is
$U(\alpha)=\exp\!\left[-\frac{i}{2}\,\delta\,\mathbf{n}(\alpha)\cdot\bm{\sigma}\right],$ (35)
and the corresponding Poincaré-sphere rotation is given by Eq. (3) with $\mathbf{n}\to\mathbf{n}(\alpha)$. The antisymmetric generator is
$A(\alpha)=\sin\delta\,[\mathbf{n}(\alpha)]_{\times}=[\bm{\Omega}(\alpha)]_{\times},\qquad\bm{\Omega}(\alpha)=\sin\delta\,\mathbf{n}(\alpha).$ (36)
Let the input state be a linear polarization along the $S_{1}$ axis,
$\mathbf{s}_{\mathrm{in}}=(1,0,0)^{\mathsf{T}}.$ (37)
As the plate rotates quasistatically from $\alpha=0$ to $\alpha=2\pi$ the Stokes vector $\mathbf{s}(\alpha)$ traces a closed curve $C$ on the Poincaré sphere. For a quarter-wave plate ($\delta=\pi/2$) this curve is a circle at a fixed latitude $\theta$ determined by the relative orientation between the input polarization and the eigenstates of the plate, while for a half-wave plate ($\delta=\pi$) it spans a larger portion of the sphere. In all cases, the instantaneous motion of $\mathbf{s}(\alpha)$ is governed by
$\frac{d\mathbf{s}(\alpha)}{d\alpha}=\bm{\Omega}(\alpha)\times\mathbf{s}(\alpha),$ (38)
and the geometric phase accumulated over a full $2\pi$ rotation is given by
$\gamma_{\mathrm{geom}}(C)=-\frac{1}{2}\,\Omega_{\mathrm{solid}}(C),$ (39)
where $\Omega_{\mathrm{solid}}(C)$ is the solid angle enclosed by the trajectory.
The explicit evaluation of $\Omega_{\mathrm{solid}}(C)$ depends on $\delta$ and on the initial state $\mathbf{s}_{\mathrm{in}}$, but its origin can be traced entirely to the antisymmetric generator $A(\alpha)$, since the symmetric part $S(\alpha)$ of the Mueller block does not enter the kinematic equation. For example, for a quarter-wave plate and a suitably chosen input state, the trajectory is a circle of colatitude $\theta$ on the sphere, and one obtains
$\gamma_{\mathrm{geom}}(C)=-\pi\bigl(1-\cos\theta\bigr),$ (40)
in agreement with the standard PB-phase result for such configurations. The present formalism shows that this phase can be read directly from $\bm{\Omega}(\alpha)$, which is determined by the antisymmetric part of the Mueller matrix of the rotating plate.
More generally, the optical retarding behavior of an arbitrary linear polarimetric transformation described by a Mueller matrix $M$ is completely captured by a pair of equivalent ideal retarders (the entrance and exit retarders), which encode the geometric phase associated with each pair of incident and emerging polarization states [16].
V.2 Quantum example: spin-1/2 in a cyclic magnetic field
As a quantum counterpart, consider a spin-$1/2$ particle in a magnetic field of fixed magnitude $B$ whose direction precesses around the $z$ axis at a fixed polar angle $\theta$. The Hamiltonian is
$H(t)$ $=-\frac{\hbar\omega}{2}\,\mathbf{n}(t)\cdot\bm{\sigma},$ (41)
$\mathbf{n}(t)$ $=\bigl(\sin\theta\cos\varphi(t),\;\sin\theta\sin\varphi(t),\;\cos\theta\bigr)^{\mathsf{T}}.$
where $\omega$ is the Larmor frequency and $\varphi(t)$ increases monotonically from 0 to $2\pi$ during a period $T$. The corresponding unitary evolution operator $U(t)$ is of the form (24), and its adjoint action on the Bloch vector is described by Eq. (25) with $\mathbf{n}\to\mathbf{n}(t)$.
If the system is prepared initially in the instantaneous ground state of $H(0)$ and the field is varied adiabatically, the Bloch vector $\mathbf{r}(t)$ follows $\mathbf{n}(t)$ and traces a cone of polar angle $\theta$ on the Bloch sphere. After one cycle the state returns to itself up to a phase, and the Berry phase is [3]
$\gamma_{\mathrm{geom}}(C)=-\pi\bigl(1-\cos\theta\bigr),$ (42)
which coincides with $\frac{1}{2}$ times the solid angle $2\pi(1-\cos\theta)$ enclosed by the cone. From the viewpoint of the present formalism, this phase arises from the tangential component $\bm{\Omega}_{\perp}(t)$ of the angular velocity of $\mathbf{r}(t)$, where
$A_{U}(t)=[\bm{\Omega}(t)]_{\times},\qquad\bm{\Omega}(t)=\sin\delta(t)\,\mathbf{n}(t),$ (43)
and $\delta(t)$ is related to the accumulated dynamical angle. The symmetric part $S_{U}(t)$ of the adjoint map does not contribute to $\gamma_{\mathrm{geom}}$.
For nonadiabatic evolutions where $\mathbf{r}(t)$ does not exactly follow $\mathbf{n}(t)$, the trajectory $C$ on the Bloch sphere changes but the geometric phase is still given by Eq. (32), with $\bm{\Omega}(t)$ determined by the antisymmetric generator $A_{U}(t)$. The explicit value of $\gamma_{\mathrm{geom}}$ now depends on the detailed dynamics, but its geometric origin and its dependence on $\bm{\Omega}_{\perp}(t)$ remain unchanged. This illustrates the fact that the antisymmetric adjoint generator provides a unified description of geometric phases in both adiabatic and nonadiabatic regimes.
VI Operational aspects and possible extensions
The identification of the antisymmetric adjoint generator as the universal kernel of geometric phase has direct operational implications in both classical polarization and quantum process characterization.
VI.1 Extraction from measured Mueller matrices
In classical polarization optics, the Mueller matrix associated with the action of a linear system can be measured experimentally using standard polarimetric techniques. The procedure to extract the entrance and exit (ideal) retarders of a general Mueller matrix is described in Ref. [16]. For a given ideal retarder, the Mueller matrix has the block form (5), with a $3\times 3$ submatrix $R$ that must belong to $\mathrm{SO}(3)$. Once $R$ is known, its antisymmetric part
$A=\frac{1}{2}\left(R-R^{\mathsf{T}}\right)$ (44)
is directly obtained and can be written as $A=[\bm{\Omega}]_{\times}$. The vector $\bm{\Omega}$ specifies the instantaneous angular velocity of the Stokes vector under the action of the retarder and, as shown in Sec. III, fully determines the geometric phase associated with any trajectory generated by the device.
This provides a simple criterion for identifying and controlling geometric-phase effects in polarization experiments. For instance, if a composite optical system is designed such that the net antisymmetric generator $A$ vanishes over a given range of parameters, then its net geometric phase is suppressed, even if the symmetric part of the Mueller block induces nontrivial transformations of Stokes vectors. Conversely, by tailoring $A$ one can design devices that implement prescribed geometric phases with a given trajectory on the Poincaré sphere.
VI.2 Quantum process tomography and adjoint maps
In quantum information experiments, the evolution of a qubit is often characterized via quantum process to- mography [15, 17]. In the Bloch representation, a general quantum channel $\mathcal{E}$ acts as
$\rho\mapsto\mathcal{E}(\rho),\qquad\mathbf{r}_{\mathrm{out}}=R_{\mathcal{E}}\,\mathbf{r}_{\mathrm{in}}+\mathbf{t},$ (45)
where $R_{\mathcal{E}}$ is a real $3\times 3$ matrix and $\mathbf{t}$ is a translation vector. For unitary channels one has $\mathbf{t}=0$ and $R_{\mathcal{E}}\in\mathrm{SO}(3)$, so that the antisymmetric part
$A_{\mathcal{E}}=\frac{1}{2}\bigl(R_{\mathcal{E}}-R_{\mathcal{E}}^{\mathsf{T}}\bigr)=[\bm{\Omega}]_{\times}$ (46)
coincides with the antisymmetric generator $A_{U}$ of Sec. IV. Therefore, once $R_{\mathcal{E}}$ has been reconstructed from experimental data, the geometric content of a unitary evolution is contained entirely in $A_{\mathcal{E}}$.
This observation suggests using the antisymmetric part of the reconstructed adjoint map as a diagnostic tool for geometric phases in noisy or imperfect implementations of qubit gates. In scenarios where $\mathcal{E}$ is close to a unitary channel, $A_{\mathcal{E}}$ provides an effective geometric generator that can be compared with the ideal design and used to detect deviations in the implemented geometric phase.
VI.3 Extensions beyond unitary two-level dynamics
The present analysis focuses on unitary dynamics in two-level systems, where the state space of pure states is the two-sphere and the adjoint action of $\mathrm{SU}(2)$ is represented by $\mathrm{SO}(3)$ rotations. It is natural to ask how far these results can be extended.
One direction is to consider more general completely-positive trace-preserving (CPTP) maps on qubits. In this case $R_{\mathcal{E}}$ need not be orthogonal, and the antisymmetric part $A_{\mathcal{E}}$ no longer generates a pure rotation. Nevertheless, $A_{\mathcal{E}}$ continues to encode the infinitesimal antisymmetric contribution to the mapping of Bloch vectors and may serve as a useful indicator of geometric features even in the presence of decoherence. A systematic analysis of geometric phases for nonunitary evolutions in terms of $A_{\mathcal{E}}$ lies beyond the scope of this work but appears as a natural extension.
Another direction is to explore higher-dimensional systems, such as spin-$j$ particles or multilevel atoms, whose pure-state manifolds can often be embedded in higher-dimensional analogues of the Bloch sphere. In these cases the relevant Lie algebras are $\mathfrak{su}(N)$ and their adjoint representations, which contain antisymmetric generators associated with rotations in generalized Bloch spaces. It is plausible that an appropriate generalization of the present construction identifies antisymmetric adjoint generators as carriers of geometric-phase information in those settings as well.
VII Conclusions
We have shown that the antisymmetric part of the adjoint $\mathrm{SU}(2)$ generator provides a universal algebraic kernel for geometric phases in two-level systems, encompassing both classical polarization optics and quantum qubits. In the classical setting of ideal retarders, the antisymmetric $3\times 3$ block of the Mueller matrix is proportional to the cross-product matrix of a vector $\bm{\Omega}$, which governs the angular velocity of the Stokes vector on the Poincaré sphere. The Pancharatnam–Berry phase is completely determined by the tangential component of this angular velocity and can be expressed as $\tfrac{1}{2}$ times the solid angle enclosed by the trajectory of the Stokes vector. The symmetric part of the Mueller block is geometrically neutral in the sense that it does not contribute to the geometric phase.
In the quantum setting, the adjoint action of any $\mathrm{SU}(2)$ unitary operator on the Bloch sphere is described by a rotation matrix whose antisymmetric part is again of the form $\bm{\Omega}(t)]_{\times}$. The Bloch vector of a pure qubit state evolves according to a kinematic equation driven by $\bm{\Omega}(t)$, and the geometric phase associated with cyclic evolutions is given by an expression formally identical to that in the classical case. This holds independently of whether the evolution is adiabatic or nonadiabatic, and it reproduces Berry’s phase in the adiabatic limit. Geometrically, the antisymmetric adjoint generator encodes not only the magnitude of the angular velocity on the Poincaré/Bloch sphere but also its handedness, so that it carries the chirality of the evolution and fixes the sign of the geometric phase.
These results establish a direct and operational bridge between classical polarization experiments, described by Mueller matrices, and quantum two-level dynamics, described by unitary adjoint maps on the Bloch sphere. In both cases, the complete geometric content of the evolution is encoded in the antisymmetric adjoint generator, which can be extracted from measured Mueller matrices or from quantum process tomography. Beyond its unifying role, this framework suggests concrete strategies for designing, controlling, and diagnosing geometric-phase effects in optical and quantum platforms, and it points naturally towards extensions to nonunitary channels and higher-dimensional systems where antisymmetric adjoint generators may play a similar geometric role.